OpenStudy (rainbowdashrocks):

A gardener weeds 2/3 of a flower bed in 6 min. How much of the flower bed did the gardener weed in 1 min?

OpenStudy (rainbowdashrocks):

@daisyduck04

OpenStudy (boldjon):

(2/3 bed) / (6min) = ('x' bed) / (1min) Cross-multiply: (2/3 bed * 3min) = ('x' bed * 6min) Divide each side by 6min: 1/3 bed = 'x' bed

OpenStudy (rainbowdashrocks):

ok